OpenStudy (studygurl14):

check answer @jim_thompson5910

OpenStudy (studygurl14):

420

rishavraj (rishavraj):

nope.....

OpenStudy (studygurl14):

aw man. :(

OpenStudy (studygurl14):

i did it again and got the same answer...?

Miracrown (miracrown):

Use the formula for arithmetic series

OpenStudy (studygurl14):

wait, no i figured out where i was making the mistake.

OpenStudy (studygurl14):

168

rishavraj (rishavraj):

just plug values in equation of finding the sum of a AP S = \[\frac{ n }{ 2 }[2a_1 + (n - 1)d]\] just plug the values

jimthompson5910 (jim_thompson5910):

Sum of the first n terms of an arithmetic sequence \[\Large S_n = \frac{n*(a_1 + a_n)}{2}\] \(\Large a_1\) is the first term \(\Large a_n\) is the nth term \(\Large a_n = a_1 + d(n-1)\) where d is the common difference

jimthompson5910 (jim_thompson5910):

The formula I posted leads to what rishavraj posted after you plug in \(\Large a_n = a_1 + d(n-1)\) and simplify a bit

Miracrown (miracrown):

\(\color{blue}{\text{Originally Posted by}}\) @StudyGurl14 168 \(\color{blue}{\text{End of Quote}}\) Nope

OpenStudy (studygurl14):

....darn

Miracrown (miracrown):

The answer is less important than understanding the method

OpenStudy (studygurl14):

10+(17-1)(4) = 74 \(\large\frac{4(10+74)}{2}=168\) @Miracrown

jimthompson5910 (jim_thompson5910):

that '4' should be '17' (since it's the value of n)

rishavraj (rishavraj):

@StudyGurl14 its \[\frac{ 17 }{ 2 }[20 + (16)4]\]

OpenStudy (studygurl14):

oh, oops. ...lol....

Miracrown (miracrown):

What they said

OpenStudy (studygurl14):

4 is the rate, not the term. my mistake

Miracrown (miracrown):

you confused n with d

OpenStudy (studygurl14):

714?

Miracrown (miracrown):

Yesh

OpenStudy (studygurl14):

thanks everyone.